If C = [10m, 30 degrees] and D =[25m, 130 degrees], whatt is the direction of the sum of these two vectors

Grade:12

1 Answers

Arun
6 years ago
The angle between the given vectors is 130-30 = 100 degree. From the Law of cosines the magnitude is
R2 = 102 + 252 - 2*10*25*cos (100) = 811.82 m2
R = 28.5 m
Using the law of Sinuses we can find the angle between R and C:
28.5/sin100 = 25/sinx , sinx = 0.86, x = 60 degrees.
Thus the direction of the vector R is 180-60 + 30 = 150 degree.
Answer: 28.5 m, 150 degree.